let a and b be two non null vectors such that |a+b|=|a-2b| . Then the value of |a|/|b| May be : 

 

(a) 1/4  (b) 1/8  (c) 1 (d) 2.  Pls solve the question and also give the reason for every step

Grade:11

1 Answers

Aksh
6 years ago
|a+b|=|a-2b|By vector law√a*a+b*b+2abcostheta=√a*a+4b*b-4abcostheta By squaring both sides a*a+b*b+2abcostheta=a*a+4b*b-4abcostheta6abcostheta=3b*b2a/b=1/costhetaa/b=1/2costhetsOnly theta=60 can give answer in options givenThus a/b=1/2*1/2=1Hence a/b may be 1i.e option(c).